# Building Confidence in Deep Generative Protein Design

### v1

@@ 3.5 Structural Phylogenetics
Two methods were used to calculate pairwise structural distances between generated and experimental structures, creating a distance matrix for constructing structural phylogenetic trees (Figure 7). After normalizing branch lengths, we generated a summary tree using SumTrees (Moreno et al., 2024), shown in Figure 2B. To ensure reliable $Q_{\text{score}}$ calculations, residue count differences between compared structures were limited to within 10% (Malik et al., 2020). Thus, we included proteins with residue counts of 243-297 for $\beta$-lactamases, 102-124 for cytochrome c, 216-261 for GFP, and 166-187 for Ras proteins, excluding class C $\beta$-lactamases ($\sim 360$ residues) from the tree but retaining them for model training and analysis.
The summary tree illustrates the potential of structural phylogenies, with each protein family clustering on distinct branches and subtypes (such as metallo-$\beta$-lactamases) grouping closely. Generated structures mostly integrate within small branches of experimental proteins, suggesting a close structural relationship. However, the models may not fully capture the specificity of class D $\beta$-lactamases, as no generated samples clustered in this class.
SM samples form dense clusters, while FM samples show greater diversity, positioning closer to the tree root but near experimental proteins without creating new branches, indicating higher novelty in FM outputs.
3.6 Molecular Dynamics
Figure 3: Stability assessment of MD simulations across proteins using various metrics. (A, C-E) Distributions of (A) RMSD, (C) radius of gyration ($R_{g}$), (D) potential energy, and (E) secondary structure counts throughout the simulation. Interquartile ranges and whiskers show metric variation; high-quality structures have medians close to experimental values with narrow ranges. (B) Residue fluctuation (RMSF) during simulation, where stable structures exhibit lower RMSF values and trends similar to the experimental structure.
Structures that appear reasonable may, in fact, be unstable due to molecular dynamics, water interactions, and entropy, which are not considered during generation. To evaluate the dynamic stability of the generated structures, we conducted MD simulations under physiological conditions and analyzed their time-dependent behavior.
Homology modeling of side-chains.
We used homology modeling to add side-chains to the protein backbone. After determining the optimal sequence (Section 3.4), we selected template proteins with at least 45% sequence identity and a $\text{TM}_{\text{score}}$ of 0.75 or higher using FoldSeek (van Kempen et al., 2023). Sequences were aligned with Clustal Omega, and MODELLER (Šali & Blundell, 1993) generated possible side-chain conformations using statistical potentials and rotamer libraries, with backbone fixed. We chose the final side-chain arrangement based on lowest energy and minimal steric clashes, verifying model quality with PROCHECK (Laskowski et al., 1993, 1996) and WHAT_CHECK (Hooft et al., 1996), discarding low-quality models.
Simulation.
After adding hydrogen atoms with Reduce2 from the computational crystallography toolbox (Grosse-Kunstleve et al., 2002), we conducted MD simulations with GROMACS (Abraham et al., 2015). The protein was placed in an octahedral box with at least 15 Å̊$\mathring{\mathrm{A}}$ from the edges, using the CHARMM36 force field (July 2022) (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Yu et al., 2012; Soteras Gutiérrez et al., 2016) for intermolecular interactions. Following vacuum energy minimization, the system was solvated, neutralized to 150 mM Na+ and Cl-, and minimized again. The system was then heated to 310 K under NVT [^2] Constant number of particles, volume, and temperature. conditions and equilibrated at 1 atm under NPT [^3] Constant number of particles, pressure, and temperature. conditions, followed by a 10 ns production run. Details are provided in Appendix F.
Stability.
Dynamic stability were analyzed as follows:
@@ 3.7 Protein-ligand Docking
Figure 4: Comparison of experimentally determined bindings (bold boxes) with predictions from blind docking simulations, using models closest to experimental data. Ligands are shown in green (experimental data), cyan (predictions from experimental receptors), magenta (predictions from SM samples), and yellow (predictions from FM samples). $\Delta G$ is in kcal/mol, with RMSD indicating deviation from experimental ligand positions (in Å). Dockings labeled “MD” represent protein-ligand complex MD simulations after blind docking.
Using AutoDock Vina (Trott & Olson, 2009; Eberhardt et al., 2021), we predicted optimal binding modes between generated structures and their family-specific ligands. We performed blind docking, scanning the entire protein surface for potential ligand binding sites without prior knowledge of binding pockets. The grid box covered the entire protein, and identical configurations were applied to both generated and experimental structures to evaluate if deviations in generated samples fell within an acceptable range, thus assessing their functional viability. Details on simulation settings are provided in Appendix G.
[…]
@@
Mutations at Glu166 and Asn170 in class A $\beta$-lactamases (Figure 1C) can form a stable acyl-enzyme intermediate, disrupting deacylation (Chen & Herzberg, 2001). Only asparagine and glycine at position 170 preserve WT-like function (Brown et al., 2009), and this conservation is retained in generated structures like FM-4 (Figure 4A), suggesting they may retain the ability to inactivate penicillin antibiotics.
@@
The generated cytochrome c-like structures retain conserved residues found in the WT (Figure 1D; 1HRC). SM-1 (Figure 4B) includes phosphorylatable residues Tyr58, Thr59, and Tyr107, as well as lysine residues Lys82, Lys83, and Lys96 involved in phospholipid binding. Asn80 and these lysines form an ATP-binding pocket-like structure. FM-0 shows heme iron coordinated by two cysteines, which may form stronger covalent bonds, potentially affecting electron transfer efficiency.
@@
GNP, a non-hydrolyzable GTP analog, is commonly used to simulate the GTP-bound active state of Ras proteins (Pantsar, 2020). In WT KRas bound to GNP (Figure 4D; 5UFE), residues from the Switch II region fill the pocket. Removal of the $\gamma$-phosphate relaxes the protein conformation to the GDP-bound state, opening the switch II region (Figure 4C; 4OBE) (Kauke et al., 2017).
After blind docking the generated KRas-like structures with GNP/GDP, we analyzed the complexes using protein–ligand MD simulations. Allowing flexibility in both the backbone and side-chains enabled us to capture the dynamic conformational changes, especially in the switch regions, between active/inactive states. FM-generated KRas-like structures show greater flexibility. In FM-13 (Figure 4D), GNP binding causes switch II residues to fill the pocket, similar to 5UFE. In the GDP-bound state, FM-13 adopts an open conformation, with the channel between the switches and P-loop opening, as seen in 4OBE (Figure 4C). In contrast, SM-27 are more rigid, with fewer conformational changes between GDP- and GTP-bound states.
@@ 4 Discussion
This study demonstrates the potential of SM and FM methods for generating protein backbones. After determining the sequence, we performed side-chain homology modeling. Despite sequence dissimilarities, key residues are conserved, and the generated structures are highly similar. SM better captures conserved regions, producing more rigid structures, while FM offers greater flexibility. Structural phylogenetic analysis revealed evolutionary relationships, and MD and docking simulations confirmed stability and functionality. The presented pipeline and evaluation protocols build confidence for broader applications in protein generation.
@@
These generative processes differ from protein synthesis simulations and do not reflect human hierarchical understanding. Limited interpretability reduces their utility for further downstream analyses and industrial applications. Although generated proteins may resemble real ones and include similar conserved residues and conformational changes, small variations could significantly impact function. Simply aiming for diversity and novelty does not guarantee quality.
Generative models have also shown great potential as efficient surrogates for MD simulations, which are often computationally intensive for complex systems or long-term behaviors. Jing et al. (2024) developed a framework to simulate molecular trajectories using generative models. Similarly, Viguera Diez et al. (2023) used generative models to enhance sampling of slow degrees of freedom, covering the sample space more effectively than traditional MD.
@@
Acknowledgments
The authors would like to thank Charlie Harris and Gos Micklem for their fruitful discussions and valuable feedback on this work.
Appendix A Protein Backbone Generation
A.1 $\mathrm{SE}(3)$ Decomposition into $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$
The Special Euclidean group $\mathrm{SE}(3)$ describes the rotations and translations in 3D space. An element of the $\mathrm{SE}(3)$ can be represented by a $4\times 4$ matrix:
@@ A.1 $\mathrm{SE}(3)$ Decomposition into $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$
A.2 Protein Backbone Representations
@@
Following the work of Yim et al. (2023) and Bose et al. (2023), we adopt the more compact backbone rigid groups from AlphaFold (Jumper et al., 2021) to represent protein backbone structures in 3D space. A backbone rigid group consists of the main chain atoms (N, Cα, C, O) within a single residue (Figure 1A), where the their geometric relationships (relative positions and orientations) are highly stable. The position and orientation of the group is transformed as a whole, without accounting for individual atomic movements, simplifying the computation and reducing structural errors caused by excessive degrees of freedom.
Assuming experimentally measured ideal chemical bond angles and lengths, models learn how the rigid transformation (or frame) $\bm{\mathsf{T}}_{i}$ of each residue $i\in[1,N]$ acts on idealized coordinates $[\mathrm{N}^{\star},\mathrm{C}_{\alpha}^{\star},\mathrm{C}^{\star}]\in\mathbb{R}^{3}$ (centered at $\mathrm{C}_{\alpha}^{\star}$), so that the transformed coordinates match the actual coordinates as closely as possible:
@@
A.3 $\mathrm{SE}(3)$ Score Matching
@@
Instead, score-based models estimates tractable conditional score $\nabla\log p_{t|0}$ through SM (Vincent, 2011), using a neural network $s(\theta,t,\cdot)$ trained by minimizing both [^4] One adds weights $1/\mathbb{E}[\lVert\nabla_{\mathbf{R}}\log p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})\rVert^{2}]$ to Equation 7 and $(1-e^{-t})/e^{-t/2}$ to Equation 8 for simplicity.:
@@
A.4 $\mathrm{SE}(3)$ Flow Matching
@@ A.5 $\mathrm{SE}(3)$ Invariance
In the context of SM, to ensure translation invariance on $\mathbb{R}^{3}$, one apply a projection matrix $\mathbf{P}\in\mathbb{R}^{3N\times 3N}$ that removes the center of mass $\frac{1}{N}\sum_{i=1}^{N}\bm{\mathsf{x}}_{i}$. It results in an invariant measure on $\mathrm{SE}(3)^{N}$, denoted as $\mathrm{SE}(3)^{N}_{0}$. Since the Brownian motion on $\mathrm{SO}(3)$ and the score $\nabla_{\bm{\mathsf{R}}}\log p_{T-t}$ are both rotation-invariant, Equation 5 is $\mathrm{SO}(3)$-invariant. Consequently, Yim et al. (2023) derive the following $\mathrm{SE}(3)$-invariant forward process
@@ A.7 Model Architecture
The networks $s(\theta,t,\cdot)$ involved in SM and $v(\theta,t,\cdot)$ involved in FM models, as reviewed in Section 2, can share a common high-level architecture (Yim et al., 2023; Bose et al., 2023; Anand & Achim, 2022).
@@
Given node embedding dimensions $D_{h}$ and edge embedding dimensions $D_{e}$, node embeddings $\mathbf{h}\in\mathbb{R}^{N\times D_{h}}$ are derived from residue indices $\textbf{i}=\{1,\dots,N\}$ and time-step information $\textbf{t}=\{0,\Delta t,\dots,T\}$, while edge embeddings $\mathbf{E}=\in\mathbb{R}^{N\times N\times D_{e}}$ integrate additional features, such as relative sequence distances $j-i$ for any $i,j\in[1,N]$ (Figure 5A). Self-conditioning on the predicted $\mathrm{C}_{\alpha}$ displacements is also applied:
@@
where $\bm{\mathsf{x}}^{*}$ denotes the coordinates for $\mathrm{C}_{\alpha}$ predicted through self-conditioning, and $v_{1},\dots,v_{B}$ are bins spaced uniformly from 0 and $B$ angstroms. These initial features are encoded using multilayer perceptrons (MLPs) along with sinusoidal embeddings (Vaswani et al., 2017).
@@
Appendix B Amino Acid Count Distributions in Experimental Proteins Used for Training
Figure 6: (A-D) Distribution of amino acid sequence lengths (aa) in experimentally determined protein structures used for training.
Appendix C Structural Phylogenetic Trees
Figure 7 shows structural phylogenetic trees generated using $Q_{\text{score}}$ and the 3Di alphabet. Both approaches reveal evolutionary relationships between the generated and experimental structures. However, compared to the $Q_{\text{score}}$-based tree, the 3Di alphabet tree shows closer and less distinguishable evolutionary relationships among samples from different protein families, especially between the GFP and cytochrome c samples.
Figure 7: Structural phylogenetic trees constructed using (A) $Q_{\text{score}}$ and (B) 3Di alphabet. Members of different protein families are grouped into semi-circular rings. The colored branches and nodes represent the generated protein structures.
Appendix D Conserved Residue Consistency
Figure 8: Normalized evolutionary rates were mapped onto sequences, with black indicating rapidly evolving positions and orange indicating conserved positions. The heat map displays each column as the optimal predicted sequence of a generated sample, with rows representing specific positions in these sequences. The bar chart (exp. identity) shows the average pairwise sequence identity at each position from the MSAs of experimental proteins. Taller bars signify positions where most experimental protein share the same residue. Consistency between generated and experimental sequences at conserved positions appears as uniform orange coloring in the heat map, aligning with the taller bars.
Appendix E Evaluation of Side-Chain Homology Modeling

### v2

@@ 3.5 Structural Phylogenetic Tree
Pairwise structural distances between generated and experimental structures were computed using two methods, forming a matrix for phylogenetic tree construction. After normalizing branch lengths, a summary tree was generated with SumTrees (Moreno et al., 2024) (Figure 9). The topology shows that both the generated and the experimentally derived proteins cluster strongly by family, without intermixing among these groups.
Structure-informed tree outperforms sequence-only tree.
Sequence-based phylogenetic trees were constructed by aligning sequences with Clustal Omega and inferring trees with FastTree (Price et al., 2009). Structure trees better preserve natural taxonomic groupings than sequence trees (Figure 10AB), especially at moderate to low sequence identity, and they achieve higher Taxonomic Congruence Scores (Tan et al., 2015) (Appendix G, Figure 10D), indicating closer alignment with the known taxonomy (Appendix G).
3.6 Dynamic Stability
@@ 3.7 Protein-ligand Docking
Mutations at Glu166 and Asn170 in class A $\beta$-lactamases (Figure 4A) can form a stable acyl-enzyme intermediate, disrupting deacylation (Chen & Herzberg, 2001). Only Asn and Gly at 170 preserve WT-like function (Brown et al., 2009), and this conservation is retained in generated samples like FM-4 (Figure 12A), suggesting they may retain the ability to inactivate penicillin antibiotics. Notably, Asn/Gly at position 170 occurred in 23/50 FM samples and 18/50 SM samples, and when extended to positions 169 to 171, in 30/50 FM samples and 22/50 SM samples, rates higher than expected by chance.
@@
The generated cytochrome c-like structures retain conserved residues found in the WT (Figure 4B; 1HRC). SM-1 (Figure 12B) includes phosphorylatable residues Tyr58, Thr59, and Tyr107, as well as lysine residues Lys82, Lys83, and Lys96 involved in phospholipid binding. Asn80 and these lysines form an ATP-binding pocket-like structure. FM-0 shows heme iron coordinated by two cysteines, which may form stronger covalent bonds, potentially affecting electron transfer efficiency.
@@
GNP, a non-hydrolyzable GTP analog, is commonly used to simulate the GTP-bound active state of Ras proteins (Pantsar, 2020). In WT KRas bound to GNP (Figure 12D; 5UFE), the Switch II region residues fill the pocket. Removing of the $\gamma$-phosphate relaxes the conformation to the GDP-bound state, opening the switch II region (Figure 12C; 4OBE) (Kauke et al., 2017).
After blind docking the generated KRas-like structures with GNP/GDP, we analyzed the complexes using protein–ligand MD simulations. Allowing flexibility in both the backbone and side-chains enabled us to capture the dynamic conformational changes, especially the WT-like transitions in the switch regions, between active/inactive states. FM-generated structures show greater flexibility. In FM-13 (Figure 12D), GNP binding causes switch II residues to fill the pocket, similar to 5UFE. In the GDP-bound state, FM-13 adopts an open conformation, with the channel between the switches and P-loop opening, as seen in 4OBE (Figure 12C). In contrast, SM-27 are more rigid, with fewer conformational changes between GDP- and GTP-bound states.
@@ 4 Discussion
The generated samples for the most part capture the observed evolutionary diversity. However, there are regions of the observed diversity that are not so well covered (Figure 9) (such as GFP and class A $\beta$-lactamases), raising concerns about potential overfitting and limited generalizability. Common metrics are inadequate for quantifying overfitting in generative models (Arora et al., 2017), especially given the variability in sample sizes and quality across families.
5 Guidelines
Here, we outline a set of guidelines and best practices for designing and validating proteins using deep generative methods: (1) Implement physical constraints during the generation process to produce realistic structures without severe steric clashes. (2) Generate multiple candidate sequences for each design and prioritize those most likely to fold into the target structure. (3) Integrate sequence design and backbone validation early to ensure compatibility between the sequence and the backbone model. (4) Identify and preserve conserved residues in the designs to maintain critical functional motifs and overall structural integrity. (5) Construct all-atom models of the designs and refine them to relieve any strain or clashes introduced during initial modeling. (6) Compare the designs to known protein family structures or fold templates to ensure consistency with known protein folds. (7) Perform molecular dynamics simulations to verify that the proposed structures remain stable under physiologically relevant conditions. (8) Perform docking analyses to evaluate the viability of proposed ligand-binding sites. (9) Experimental validation is always essential to confirm the intended function and structural integrity of each candidate. (10) See Appendix J for circumstances under which implementation of these guidelines can be challenging.
6 Conclusion
This study shows the potential of methods based on deep generative models for designing proteins with high structural fidelity and functional plausibility. SM better captures conserved regions, producing more rigid structures, while FM offers greater flexibility (Figure 8). Although the optimized sequences for the generated structures exhibit some differences from natural family members, they preserve many functionally essential conserved residues (Figure 7). In structural phylogenetic trees (Figure 9), these designs cluster with their respective families, suggesting that their geometries could capture some family‐specific signatures indicative of common ancestry or function. MD and docking simulations are consistent with the designs remain properly folded and stable under physiological conditions over time (Figure 11), and can accommodate their family-specific ligands in binding pockets that closely correspond to known active sites (Figure 12). Some designs undergo conformational changes upon binding that mirror the allosteric behavior observed in WT (Figure 12CD).
@@ 6 Conclusion
Acknowledgements
The authors would like to thank Charlie Harris, Simon Mathis and Stephen Eglen for their fruitful discussions and valuable feedback on this work.
Appendix A Deep Generative Protein Design Workflow
Figure 1: A schematic overview of the deep generative protein design pipeline and its evaluation protocols.
Appendix B Protein Backbone Generation
B.1 $\mathrm{SE}(3)$ Decomposition into $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$
The Special Euclidean group $\mathrm{SE}(3)$ describes the rotations and translations in 3D space. An element of $\mathrm{SE}(3)$ can be represented by a $4\times 4$ matrix:
@@ B.1 $\mathrm{SE}(3)$ Decomposition into $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$
B.2 Protein Backbone Representations
Figure 2: Protein backbone with dihedral angles $\psi$ and $\phi$.
@@ B.2 Protein Backbone Representations
Following the work of Yim et al. (2023) and Bose et al. (2023), we adopt the more compact backbone rigid groups from AlphaFold (Jumper et al., 2021) to represent protein backbone structures in 3D space. A backbone rigid group consists of the main chain atoms (N, Cα, C, O) within a single residue (Figure 2), where their geometric relationships (relative positions and orientations) are highly stable. The position and orientation of the group is transformed as a whole, without accounting for individual atomic movements, simplifying the computation and reducing structural errors caused by excessive degrees of freedom.
Assuming experimentally derived ideal chemical bond angles and lengths, models learn how the rigid transformation (or frame) $\bm{\mathsf{T}}_{i}$ of each residue $i\in[1,N]$ acts on idealized coordinates $[\mathrm{N}^{\star},\mathrm{C}_{\alpha}^{\star},\mathrm{C}^{\star}]\in\mathbb{R}^{3}$ (centered at $\mathrm{C}_{\alpha}^{\star}$), so that the transformed coordinates match the actual coordinates as closely as possible:
@@
B.3 $\mathrm{SE}(3)$ Score Matching
@@
Instead, score-based models estimate tractable conditional score $\nabla\log p_{t|0}$ through SM (Vincent, 2011), using a neural network $s(\theta,t,\cdot)$ trained by minimizing both [^3] One adds weights $1/\mathbb{E}[\lVert\nabla_{\mathbf{R}}\log p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})\rVert^{2}]$ to Equation 6 and $(1-e^{-t})/e^{-t/2}$ to Equation 7 for simplicity.:
@@
B.4 $\mathrm{SE}(3)$ Flow Matching
@@ B.5 $\mathrm{SE}(3)$ Invariance
In the context of SM, to ensure translation invariance on $\mathbb{R}^{3}$, one applies a projection matrix $\mathbf{P}\in\mathbb{R}^{3N\times 3N}$ that removes the center of mass $\frac{1}{N}\sum_{i=1}^{N}\bm{\mathsf{x}}_{i}$. It results in an invariant measure on $\mathrm{SE}(3)^{N}$, denoted as $\mathrm{SE}(3)^{N}_{0}$. Since the Brownian motion on $\mathrm{SO}(3)$ and the score $\nabla_{\bm{\mathsf{R}}}\log p_{T-t}$ are both rotation-invariant, Equation 4 is $\mathrm{SO}(3)$-invariant. Consequently, Yim et al. (2023) derive the following $\mathrm{SE}(3)$-invariant forward process
@@ B.7 Model Architecture
The networks $s(\theta,t,\cdot)$ involved in SM and $v(\theta,t,\cdot)$ involved in FM models, as reviewed in Appendix B, can share a common high-level architecture (Yim et al., 2023; Bose et al., 2023; Anand & Achim, 2022).
@@
Given node embedding dimensions $D_{h}$ and edge embedding dimensions $D_{e}$, node embeddings $\mathbf{h}\in\mathbb{R}^{N\times D_{h}}$ are derived from residue indices $\textbf{i}=\{1,\dots,N\}$ and time-step information $\textbf{t}=\{0,\Delta t,\dots,T\}$, while edge embeddings $\mathbf{E}\in\mathbb{R}^{N\times N\times D_{e}}$ integrate additional features, such as relative sequence distances $j-i$ for any $i,j\in[1,N]$ (Figure 3A). Self-conditioning on the predicted $\mathrm{C}_{\alpha}$ displacements is also applied:
@@
where $\bm{\mathsf{x}}^{*}$ denotes the coordinates for $\mathrm{C}_{\alpha}$ predicted through self-conditioning, and $v_{1},\dots,v_{B}$ are bins spaced uniformly from 0 to $B$ angstroms. These initial features are encoded using multilayer perceptrons (MLPs) along with sinusoidal embeddings (Vaswani et al., 2017).
@@
Appendix C Protein Families Involved in This Study
Figure 4: (A) Overlay of WT E. coli TEM1 (PDB: 1BTL; Jelsch et al. (1993); green) and its E166N acylated intermediate (PDB: 1FQG; Brown et al. (2009); white) with penicillin (PNM). (B) WT E. caballus heart cytochrome c (PDB: 1HRC; (Dickerson et al., 1967)) with heme C (HEC). (C) A. victoria GFP and its chromophore (PDB: 4KW4; Barnard et al. (2014)). (D) GDP-bound H. sapiens KRas protein (PDB: 4OBE; (Hunter et al., 2014)).
C.1 $\beta$-lactamases
$\beta$-lactamases are enzymes that deactivate $\beta$-lactam antibiotics by hydrolyzing their $\beta$-lactam ring, contributing significantly to bacterial resistance (Lee et al., 2016). Inhibiting these enzymes can restore antibiotic efficacy (Behzadi et al., 2020). Rapid diversification driven by the evolution of bacterial resistance makes $\beta$-lactamases ideal targets for protein modeling studies. For this study, we gathered structural data on 1,578 unique monomeric $\beta$-lactamases and their variants across Ambler classes (A, B, C, and D) from the BLDB (Naas et al., 2017) and the Protein Data Bank (PDB) (Berman, 2000).
Class A $\beta$-lactamases are the most prevalent, with conserved active-site residues Ser70, Glu166, and Asn170 coordinating the hydrolytic water for deacylation (Tooke et al., 2019; Brown et al., 2009). Figure 4A shows the WT $\beta$-lactamase TEM1 (1BTL) alongside its acylated E166N intermediate (1FQG).
C.2 Cytochrome c
Figure 4B shows horse cytochrome c, where a hydrophobic shell surrounds the heme group, with only $\sim 7.5\%$ of the surface available for electron transfer (Bushnell et al., 1990). The hydrophobic environment and iron coordination by His18 and Met80 maintain a high redox potential ( 260 mV) (Salemme, 1977). Phosphorylation occurs at Thr28, Thr47, Tyr48, and Tyr97 (Hüttemann et al., 2011), while Lys72, Lys73, and Lys87 bind phospholipids (Kagan et al., 2009). The ATP-binding pocket involves Glu69, Asn70, Lys88, and Lys72, Lys86, Lys87 (McIntosh et al., 1996).
C.3 Green fluorescent proteins (GFP)
GFP, first isolated from Aequorea victoria, fluoresces green when stimulated by specific wavelengths. Its core structure is an 11-strand $\beta$-barrel enclosing the chromophore (Figure 4C) (Remington, 2011). Various mutants have been engineered to enhance or modify its properties, including enhanced GFP (Cormack et al., 1996), superfolder GFP (Pédelacq et al., 2005), and color variants like YFP (Ormö et al., 1996) and BFP (Glaser et al., 2016). We collected structural data for 448 GFPs and variants from the PDB to explore this diversity.
C.4 Ras Proteins
Ras proteins, a subgroup of the small GTPase superfamily, act as molecular switches, cycling between GTP-bound (active) and GDP-bound (inactive) states to regulate cell proliferation, differentiation, migration, and apoptosis (Ladygina et al., 2011; Simanshu et al., 2017; Weinmann & Ottow, 2007). They play a key role in signaling from the cell surface to downstream pathways. Mutations that keep Ras proteins in an active state drive excessive cell growth and malignancy, making Ras inhibition a promising cancer treatment strategy (Singh & Lingham, 2002). We focused on the most common cancer-related Ras proteins (HRas, KRas, and NRas) (Cox, 2002) and obtained 511 experimentally derived structures from the PDB.
In human KRas (Figure 4D), the switch I and II regions form the key interface for effector and regulator binding (Pantsar, 2020). These regions are highly flexible, with conformations depending on GTP or GDP binding. Cancer-related mutations frequently occur in the P-loop and switch II (Pantsar, 2020).
C.5 Sequence Length Distributions in Experimentally Derived and Generated Protein Structures
Figure 5: Distributions of amino acid sequence lengths (aa) for the experimentally derived protein structures used for training (top row; orange) and the 50 backbone structures generated by each model (bottom row; blue). Target sequence lengths for the generated structures were sampled from the training data distribution, shown here for (A) $\beta$-lactamase, (B) cytochrome c, (C) GFP, and (D) Ras.
Appendix D Dihedral Angles $\psi$ and $\phi$ Distributions
Figure 6: Ramachandran plots comparing dihedral angles $\psi$ and $\phi$ distributions for generated versus experimentally derived proteins; inset shows favored (light) and disallowed (dark) regions.
Appendix E Conserved Residue Consistency
E.1 Evolutionary Rates Mapped onto Structures
Figure 7: Normalized evolutionary rates mapped onto structures, with white for rapidly evolving positions and black for conserved ones. Experimentally derived structures are highlighted with bold outlines. Refer to Figure 4 and Figure 12 for the relevant key residues and the locations of the ligand-binding pockets (represented by PNM, HMC and GDP).
E.2 Evolutionary Rates Mapped onto Sequences
Figure 8 shows that, except for $\beta$‑lactamase, the FM-generated sequences show a slightly higher average pairwise distance than those of the SM, indicating greater diversity. One possible explanation is that the experimentally derived $\beta$‑lactamase structures used for training (from four distinct Ambler classes) have more variability than the other three protein families (as also reflected in Figure 9) and that SM is more sensitive to structural variability.
Figure 8: Multiple sequence alignment of generated samples and experimentally derived reference proteins (1FQG, 1HRC, 4XWA, 4O8E) using Clustal Omega, with normalized evolutionary rates overlaid. Rapidly evolving positions are highlighted in light gray, conserved regions in black, and alignment gaps in white. In addition, a comparative analysis of variability is presented: average pairwise distances, derived from the Clustal Omega distance matrix, denote greater variability when larger values are observed.
Appendix F Summary Structural Phylogenetic Tree
Figure 9: Summary structural phylogenetic tree constructed using the $Q_{\text{score}}$ and the 3Di alphabet. Branch colors distinguish families, with orange and blue nodes representing SM- and FM-generated structures, respectively. In $\beta$-lactamases, distinct Ambler classes are differentiated by unique background colors.
Appendix G Structure-informed Trees versus Sequence-based Phylogenetic Trees
Retrieving known taxonomic lineages.
Following Moi et al. (2023), we retrieved taxonomic lineages for each experimentally derived sequence and structure within every protein family using the UniProt API (Patient et al., 2008), assuming that most genes evolve in a manner that mirrors the species tree with only occasional instances of gene loss or duplication.
Figure 10: (A) Ultrametric summary structural phylogenetic tree constructed using the $Q_{\text{score}}$ and the 3Di alphabet. (B) Ultrametric sequence-based phylogenetic tree constructed using the Clustal Omega and the FastTree pipeline. Different protein families are differentiated by distinct colored branches. Nodes in orange and blue represent structures generated by SM and FM, respectively. (C) Phylogenetic trees of distinct protein families. Top: summary structural phylogenetic tree constructed using the $Q_{\text{score}}$ metric and the 3Di alphabet. Bottom: sequence-based phylogenetic tree inferred using the Clustal Omega and FastTree pipeline. (D) Taxonomic congruence score for each node in the sequence and structure trees. On average, structural trees exhibit higher taxonomic congruence than sequence-based trees.
Using $Q_{\text{score}}$ in structural phylogenetics.
For any two structures with $N_{1}$ and $N_{2}$ residues, $Q_{\text{score}}$ is computed with TM-align (Zhang, 2005) as:
[…]
Taxonomic congruence score (TCS).
Tan et al. (2015) proposed the use of the TCS to assess how well a phylogenetic tree’s topology agrees with the known taxonomy, arguing that TCS is an unbiased measure of tree quality. Typically, trees with higher average TCS (structure trees in Figure 10D) are considered to have more accurate topologies. In this study, we evaluated the congruence between the given trees and the established taxonomy lineages derived by UniProt. Subsequently, we provide a brief overview of the bottom-up TCS implementation as described by Moi et al. (2023):
For any node $x$ in the tree, $s(x)$ is the set of taxonomic lineage labels present in the subtree rooted at $x$. If $x$ is a leaf, then $s(x)$ is defined to be the set of lineage labels (its taxonomic classification) for that leaf. If $x$ is an internal node with children $\{y_{1},\cdots,y_{N}\}$, then $s(x)=\bigcap_{i=1}^{N}s(i)$.
For each node $x$, the function $C(x)$ quantifies the congruence of that node’s grouping with taxonomy: $C(x)=|s(x)|+|s(p)|$, where $p$ is the parent nodes of $x$, and ||$\cdot$|| ⋅ | denotes the size of a set.
The overall taxonomic congruence score for the entire tree is obtained by summing the contributions of all leaves. To compare congruence across trees of different sizes, the raw total score is typically normalized.
Appendix H Evaluation of Side-Chain Homology Modeling